### v2

# From Cool Demos to Production-Ready FMware: Core Challenges and a Technology Roadmap

Abstract.
The rapid expansion of foundation models (FMs), such as large language models (LLMs), has given rise to FMware—software systems that integrate FMs as core components. While building demonstration-level FMware is relatively straightforward, transitioning to production-ready systems presents numerous challenges, including reliability, high implementation costs, scalability, and compliance with privacy regulations. Our paper conducts a semi-structured thematic synthesis to identify the key challenges in productionizing FMware across diverse data sources including our own industry experience in developing FMArts—a FMware lifecycle engineering platform and integrating it into Huawei cloud, grey literature, academic publications, hands-on involvement in the Open Platform for Enterprise AI (OPEA), organizing the AIware conference and Bootcamp, and co-leading the ISO SPDX SBOM working group on AI and datasets. We identify critical issues in FM selection, data and model alignment, prompt engineering, agent orchestration, system testing, and deployment, alongside cross-cutting concerns such as memory management, observability, and feedback integration. We discuss needed technologies and strategies to address these challenges and offer guidance on how to enable the transition from demonstration systems to scalable, production-ready FMware solutions. Our findings underscore the importance of continued research and multi-industry collaboration to advance the development of production-ready FMware.
AI-powered Software, Production-ready FMware, FMware, Productionization
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligenceccs: Software and its engineering
1. Introduction
FMware refers to software integrating foundation models (FMs), like Large Language Models (LLMs), as core components (Hassan et al., 2024b). Since ChatGPT’s release in late 2022, the FMware landscape has exploded, with over 600,000 open-source models, including FMs and other AI models, now available on platforms like Hugging Face (Santa, [n. d.]). Goldman Sachs predicts FMware could boost global GDP by 7% (gol, [n. d.]).
Building impressive demos with FMs is relatively easy, but transitioning to production-ready FMware incurs significant challenges due to its inherent complexity as a compound system (Zaharia et al., 2024; Hassan et al., 2024b). Unlike static systems, FMware integrates multiple dynamic components for real-time updates, control, and adaptability (Zaharia et al., 2024; ibm, [n. d.]). This complexity is reflected in industry adoption rates: a survey of 430 technology professionals revealed that only 10% of organizations had launched FMware in production environments (int, 2023). Respondents identified reliability, high implementation costs, latency, compliance, and privacy as the top challenges preventing them from moving from demos to production-ready FMware.
The transition to production-ready FMware is formidable due to its dynamic nature, which combines classical software engineering complexities with unique issues inherent to foundation models. Production-ready FMware must continuously evolve to meet customer expectations, requiring consistent performance, reliability, feature updates, and adherence to service level agreements (SLAs), all while managing operational costs. For instance, OpenAI’s infrastructure costs for running ChatGPT in 2023 were estimated at \$700,000 per day (Abbas, 2024). Beyond these traditional challenges, FMware developers also face unique problems like handling hallucinations, increased inference costs, and orchestrating tasks across various AI components (OPEA, [n. d.]; Hassan et al., 2024b). LinkedIn’s experience illustrates the effort required to reach production readiness: they achieved 80% functionality in a month but spent four more months to complete the remaining 20%, with diminishing returns on each additional 1% improvement (lin, 2024). Similarly, Microsoft and GitHub found that testing FMware becomes prohibitively expensive as complexity scales (Parnin et al., 2023). These real-world examples underscore the need for robust, system-based approaches to build reliable, production-ready FMware solutions (Zaharia et al., 2024; OPEA, [n. d.]; Hassan et al., 2024b).
In this paper, we employ a semi-structured thematic synthesis (more details in Section 3) to systematically identify key challenges in developing production-ready FMware. We draw on insights from multi-industry collaborations, conferences, customer meetings, hands-on development of the FMware lifecycle platform (FMArts) (Hassan et al., 2024b) and literature surveys. In doing so, we provide the first comprehensive articulation of the recurrent issues across different stages of FMware development, consolidating them into challenges that hinder productionizing FMware.
This paper complements our vision for Software Engineering 3.0 (SE 3.0), where we advocate for an AI-native, intent-first approach where AI systems evolve from task-driven copilots to intelligent collaborators (Hassan et al., 2024c). Both SE 3.0 and our paper address the need to adapt software engineering practices to manage the complexity and dynamism of AI systems like FMware. By synthesizing hands-on knowledge from building production-ready FMware and interacting with industry experts, we underscore the urgency of addressing these challenges to move from “cool” demos to robust, production-ready solutions. Our study further highlights the importance of ongoing research and multi-industry collaboration as FMware continues to evolve.
This paper is organized as follows: Section 2 covers the evolution of FMware and related work. Section 3 explains our thematic analysis approach. Section 4 outlines the FMware lifecycle and spot recurring issues at each stage that affect production readiness. Section 5 presents the key challenges hindering FMware production-readiness and suggests future directions. Finally, Section 6 concludes the study.
2. Related Work
2.1. Release and Production Readiness in Software Engineering
Despite a wealth of research on software release readiness (Al Alam et al., 2014; Koprowski et al., 2014; Al Alam et al., 2017; Port and Wilf, 2013; Patel et al., 2024), there is a notable lack of studies on production readiness (Asthana and Olivieri, 2009; Cusick, 2013). Release readiness ensures that software meets functional requirements and is ready for distribution. Whereas, production readiness also guarantees reliable operation in live environments by addressing factors like latency and operational costs. Early work by Mockus (2003) explored an analogy-based prediction method to model and forecast the workflow of work items in software projects to asses release readinees of a software. Port and Wilf (2013) explored release readiness through a case study at NASA JPL, followed by several studies quantifying and predicting release readiness factors using machine learning (Al Alam et al., 2014; Koprowski et al., 2014; Al Alam et al., 2017). More recently, Patel et al. (2024) proposed a checklist to assess the release readiness of FMware.
In assessing production readiness, Asthana and Olivieri (2009) developed a metric focused on software quality and reliability, while Cusick (2013) proposed an architecture review method. More recently, Breck et al. (2017) in their seminal work, examined the production readiness of machine learning systems, proposing a rubric to reduce technical debt and evaluate readiness. Parnin et al. (2023) documented the difficulties of building a production-ready copilots, by interviewing the team that built various copilots at Microsoft. In their influential blog post, Yan et al. ([n. d.]) shed light on some of the challenges that impact the production-readiness of FMware. Recently, Nahar et al. (Nahar et al., 2024) investigated emerging challenges, particularly those related to quality assurance, faced by developers integrating foundation models (FMs) into existing software products. Their study, based on interviews with 26 developers and a survey of 332 participants at Microsoft, provides recommendations for the development and integration of FMs into software products. Unlike their work, which is limited to insights from practitioners within a single company, our study analyzes grey literature from diverse sources to identify recurrent issues affecting all stages of the FMware lifecycle.To the best of our knowledge, this is the first study to systematically identify and organize the challenges in building production-ready FMware, highlighting core challenges and offering a technology roadmap for researchers and practitioners.
2.2. Surveys in FMware research
While several surveys have examined FMware and the role of LLMs in software engineering, they predominantly focus on academic publications or practitioner interviews. Jin et al. (Jin et al., 2024) provide a comprehensive exploration of LLMs and their application as autonomous agents in software engineering. Similarly, Fan et al. (Fan et al., 2023) survey LLMs’ potential in software development activities, including debugging and refactoring, emphasizing hybrid approaches that combine classical software engineering techniques with LLMs.
Zhang et al. (Zhang et al., 2023) investigate the impact of LLMs across five phases of the software engineering lifecycle, leveraging academic datasets to highlight technical challenges such as model tuning and evaluation. Hou et al. (Hou et al., 2024) curated a systematic review of 395 research papers to understand how LLMs are used in various software engineering tasks and on which tasks LLMs have shown success. Jiang et al. (Jiang et al., 2024) focus on LLMs for code generation, presenting a taxonomy of existing approaches based on structured academic benchmarks such as HumanEval and MBPP.
Wang et al. (Wang et al., 2024a) and others emphasize the use of LLMs in software testing, particularly for unit test generation and debugging, using standard datasets like Defects4J. However, all of these studies share a common limitation: an overreliance on academic datasets and benchmarks without addressing the contributions of open-source projects, forums, and informal knowledge exchange. Different from the aforementioned studies, Among the existing studies, Chen et al. (Chen et al., 2025) provide the most relevant foundation for our work by focusing on the challenges faced by LLM application developers. They present a comprehensive empirical study, mining 29,057 posts from the OpenAI developer forum to construct a detailed taxonomy of challenges. Their taxonomy highlights critical issues such as prompt engineering, API limitations, rate constraints, and hallucination management. To validate their findings, they extend their analysis to GitHub issues for LLaMa and Gemini, demonstrating the generalizability of their taxonomy across platforms. While their study provides actionable insights for API-driven LLM development, it overlooks challenges unique to users of smaller, self-hosted FMs, agentic development, and quality assurance activities, which are more common in enterprise environments prioritizing cost, quality, trustworthiness, and data privacy.
Unlike prior studies, which primarily rely on academic publications, this paper synthesizes insights from diverse data sources, including open-source project working group meeting notes, industry reports, grey literature (including academic publications) and informal knowledge exchanges. This holistic approach allows us to uncover the practical challenges and strategies employed in real-world FMware development. To the best of our knowledge, this is the first study to integrate academic insights, community-driven innovations, and practitioner expertise to provide a comprehensive overview of the recurrent issues that hinder the productionization of FMware across its various lifecycle stages and present a technology roadmap to address them.
3. Methodology
We conducted a semi-structured thematic synthesis, inspired by thematic analysis (Butler et al., 2024; Cruzes and Dybå, 2010), to identify and organize key challenges in developing production-ready FMware. Unlike a full thematic analysis, which involves detailed iterative coding and theme validation, we focused on systematically collecting and analyzing relevant sources, identifying issues, and grouping them into themes through expert-driven discussions.
The decision to employ a semi-structured thematic synthesis was driven by several factors. First, the sheer volume and rapid growth of publications (a recent study by Fan et al. (Fan et al., 2024) finds more than 5,000 academic publications in FMware between 2017 and 2023) in the FMware domain and the sheer scope of data (including meeting notes from several open source projects and conference reports from over 3 years) made comprehensive collection and iterative coding impractical, especially given the time constraints faced by the authors as industry researchers. Second, we mitigated potential biases by triangulating insights from diverse sources, including industry discussions, conference reports, and academic papers, ensuring a balanced and comprehensive perspective. Third, this methodology emphasizes actionable insights and practical relevance, which are critical for addressing real-world challenges in productionizing FMware. While closely following key steps of thematic analysis, such as identifying and grouping issues into themes, we adapted the process by skipping exhaustive iterative coding and instead leveraging expert judgment to refine themes.
We acknowledge that this approach has limitations, such as the absence of formal coding and the potential for subjectivity. However, these trade-offs were necessary to address the vastness and complexity of the domain, allowing us to provide a structured and systematic exploration of key challenges in transitioning FMware from demos to production-ready systems.
textbfStep 1: Data Collection. We collected data from various sources to gain insights into the challenges that researchers and practitioners encounter when productionizing FMware.
  • OPEA Project Participation: As active participants, we played a significant role in the OPEA initiative (Open Platform for Enterprise AI (OPEA) (OPEA, [n. d.])). In particular, the last author leads OPEA’s research working group. OPEA is a global multi-company collaborative initiative focused on addressing the challenges of making FMware enterprise-ready. OPEA provides detailed frameworks, architectural blueprints, and a four-step assessment for evaluating FMware in terms of enterprise readiness. We used meeting notes (some of which are publicly available (atl, 2024)) from OPEA discussions, which involved 43 organizations (e.g., Intel, AMD, RedHat, Docket, JFrog, Anyscale, LlamaIndex and SAP).
  • SPDX Working Group: First author co-led SPDX community meetings (an ISO/IEC 5692:2021 Software Bill of Materials standard (SPDX, 2021)) over the past three years, contributing to discussions on the AI Bill of Materials (AI BOM), which included discussing challenges in building production-ready FMware. The meeting notes, available publicly (SPDX, [n. d.]), were incorporated into our analysis.
  • Inception and Launch of the AIware Conferences, summit and Bootcamp: Authors of this paper spearheaded the first ACM International Conference on AI-Powered Software (AIware) (aiw, 2024), co-located with FSE 2024, the FM+SE summit series (fms, 2024) and the AIware Bootcamp (Rajbahadur et al., 2024). Engaging with participants from Google, Microsoft, GitHub, and other industry and academic professionals provided valuable insights into the challenges of productionizing FMware, which we summarized in a report used in our analysis (Hassan et al., 2024a).
  • Conference and Workshop Attendance and Reports: Since late 2022, all authors have actively participated in several top conferences, workshops, and developer meetings relevant to FMware, such as ICSE, FSE, FM+SE 2030, FM+SE Tokyo, SEMLA, OSS Summits, Ray summits and ai_dev Summits. Although these events covered a range of topics beyond the challenges of productionizing FMware, we gained valuable insights by listening to and interacting with researchers and practitioners involved in productionizing their own FMware. After each conference, the attending authors compiled detailed reports summarizing key discussions and relevant research on FMware. These reports formed the foundation for our thematic analysis.
  • Internal Industry Experience: Our collaborations with customers and development teams to understand FMware’s functional and non-functional requirements. Additionally, all authors contributed to developing and productionizing FMArts, a comprehensive lifecycle engineering platform for FMware (Hassan et al., 2024b). Our practical experience in creating FMware with FMArts and making it production-ready (Hassan et al., 2024b) and integrating it into Huawei products (e.g., Huawei Cloud) informed the thematic analysis.
  • In-depth Literature Review: We began with a review of grey literature, including blog posts, whitepapers, and developer forums, and noticed recurring discussions about the production readiness of FMware. Building on these insights, we conducted a focused review of academic papers to find supporting evidence. Through this activity, we provide a comprehensive analysis of the challenges in deploying FMware, grounding all issues in Section 4 with citations and minimizing subjective bias.
The collected materials were reviewed to identify recurring issues relevant to FMware production readiness. These insights formed the basis for subsequent grouping and theme development.
Step 2: Data Analysis and Grouping. After data collection, three of the four authors reviewed the materials to identify recurring issues. Instead of formal iterative coding, we relied on a collaborative and expert-driven approach to extract key challenges and categorize them. Issues were grouped into preliminary themes based on shared characteristics or relevance to specific stages of the FMware lifecycle (Figure 1). For example, issues related to Data and FM Alignment included challenges such as low data quality and insufficient data diversity. This approach emphasized practical categorization rather than exhaustive formal coding.
Step 3: Collaborative Discussion to Identify Themes. The authors collaboratively refined the initial groupings into overarching themes (which we present as challenges) through discussions informed by their domain expertise. These themes spanned multiple FMware lifecycle stages and were grounded in the collected data but did not follow the strict iterative refinement protocols of thematic analysis. Instead, the grouping process was guided by the practical relevance of the challenges and their applicability to real-world FMware development.
Step 4: Thematic Consolidation. In the final step, the authors revisited and refined the overarching themes to ensure they comprehensively represented the identified challenges. The process prioritized clarity and relevance for practitioners while acknowledging the absence of exhaustive formal coding. These themes are presented in Section 5, offering a systematic view of critical challenges in transitioning FMware from demos to production-ready systems.
4. Recurrent Issues in Productionizing FMware
In this section, we outline the stages of an FMware’s engineering lifecycle, as depicted in Figure 1, and highlight the recurrent issues observed.
Figure 1. FMware Lifecycle.
4.1. FM(s) Selection
The first step in developing FMware is to select the FM(s) to be used. There exist over 600,000 open-source models available (among FMs and the more traditional machine learning models) (Santa, [n. d.]). Typically, the number of FMs to be selected depends on the complexity and functionality required by the FMware. For example, a simple chatbot may only need one FM, while complex systems might require multiple FMs for its different functionalities.
Difficulty balancing functional requirements with performance and costs. Selecting FMs for FMware requires balancing functionality with performance, infrastructure, and costs (Kamath et al., 2024). Demos favor large general-purpose models like ChatGPT-4, neglecting latency and resource limits due to the typically small number of required API calls. LLaMa-70B, for example, needs 140GB VRAM to be hosted locally (Chavan et al., 2024). In production, smaller FMs are preferred for cost, latency, and privacy reasons (especially in the context of edge devices, such as smartphones) (Li et al., 2023; Wan et al., 2023), though they may lack the full capabilities of larger FMs (Chen and Varoquaux, 2024; Irugalbandara et al., 2024). However, it is important to note that even hosting smaller models can also increase ownership costs (Kamath et al., 2024).
4.2. Data and FM Alignment
After selecting FMs, the next crucial step is data and model alignment. This process tailors the FM(s) to meet the specific requirements and goals of a given FMware application (Xia et al., 2024).
Data alignment curates specialized data through labeling, generation, debugging, and testing, especially for enterprise cases (Liu et al., 2024b; Albalak et al., 2024). Domain-specific datasets are created to ensure that an FMware can perform well on the use-case specific data. Model alignment adjusts FM parameters using techniques like supervised fine-tuning and Reinforcement Learning with Human Feedback (RLHF) to ensure the FMware behaves as intended (Shen et al., 2023).
Low data quality. High data quality is crucial for reliable FMware, as poor data leads to untrustworthy outputs (Albalak et al., 2024).
Low domain coverage. Ensuring broad domain coverage is key for trustworthy FMware. Insufficient domain representation can result in biased outputs.
Low data efficiency. Accurately estimating inbound and outbound data volume is crucial for production-ready FMware, impacting both budget and performance (Chen et al., 2023). Underestimation leads to poor model performance, while overestimation wastes resources (Wang et al., 2022; Chen et al., 2023).
4.3. Prompting
Prompting involves designing and refining prompts to guide FMs toward desired outputs using precise instructions and in-context learning. While demos can get away with simple or non-optimized prompts, production-ready FMware demands meticulous engineering for consistent, reliable performance (Parthasarathy et al., 2024).
God prompts. “God prompts” that try to handle multiple tasks at once lead to unpredictable outputs and complicate debugging and maintenance (Yan et al., [n. d.]; Parthasarathy et al., 2024). In production, tasks should be split into smaller, focused prompts and be decoupled for better accuracy, debuggability and maintainability (similarly to code) (Yan et al., [n. d.]).
Lack of Built-in QA Checks for Prompts. Production-ready FMware demands robust quality assurance (QA) for prompts, including semantic and structural validation. Without built-in QA checks, prompts may produce erroneous or inappropriate responses, leading to potential failures and a loss of user trust (Kamoi et al., 2024; Ahmed and Devanbu, 2023; Wang et al., 2024b).
Lack of Prompt Portability Across FMs. Prompts optimized for one FM typically fail with others due to differing architectures and training data (Chang et al., 2024). This drastically limits portability and poses challenges to software evolution.
Absence of FM-Specific Optimizations. FM-specific optimizations can boost performance by up to 10% by tailoring prompts to each model’s architecture (Chen et al., 2024c). However, these enhancements create dependencies, complicating updates and maintenance (Sabbatella et al., 2024).
Complexity in determining failure rationale. Diagnosing whether failures result from prompt issues or FM shortcomings is difficult, often hindering the identification of the root cause and delaying effective resolution (Ashtari et al., 2023; Amujo and Yang, 2024).
Non-representative or insufficient in-context learning examples. Providing insufficient or non-representative examples for in-context learning degrades FM performance. For instance, Wang et al. (2023c) show that random in-context learning examples yield suboptimal outputs. Demos can use minimal examples, but production-ready FMware requires comprehensive, relevant examples for reliability.
4.4. Grounding
Grounding improves FM responses by linking them to external data for accuracy and relevance. Methods like Retrieval-Augmented Generation (RAG) achieve this by retrieving relevant information and using it to guide or inform responses. Effective grounding is crucial for reliable production-ready FMware.
Low information density in grounding data. Low-density grounding data impairs FM performance by limiting its ability to generate coherent responses (Laban et al., 2024; Wang et al., 2024c). In production, rich, well-structured data is essential for accurate outputs. Laban et al. (2024) show how poorly structured, vast information leads to failures in long-context tasks, while Wang et al. (2024c) emphasize efficient data compression and selection to maintain performance and relevance in complex queries.
Irrelevant grounding data. Incorporating irrelevant grounding data confuses FMs and reduces performance. Liu et al. (2024a) show that models struggle with long-context information, especially when crucial data is buried, while Cuconasu et al. (2024) note how noisy data worsens the focus of retrieval systems. For production environments, filtering and curating data is crucial to ensure only pertinent information is used.
Overcomplicating solutions with advanced techniques. Using complex methods when simpler ones would suffice adds unnecessary overhead. For instance, neural retrieval can raise computational costs without added benefits when a keyword search would be enough (Yan et al., [n. d.]; Sawarkar et al., 2024). In production, simplicity promotes scalability and maintainability, while demos often prioritize novelty (Yan et al., [n. d.]; Sawarkar et al., 2024).
4.5. Agent(s)
Agents in FMware are autonomous components that leverage the decision-making and reasoning abilities of FMs to achieve tasks with minimal human input. They dynamically execute actions, interact with environments, and collaborate with other agents. By combining capabilities from previous software generations. In this new generation agents autonomously learn, adapt, and perform multi-step tasks, enhancing the functionality of FMware.
God Agents. Monolithic “God agents” handling many tasks create complexity and maintenance challenges in production. In production, smaller specialized agents are preferred for modularity and reliability, offering better scalability and maintainability (Rasal and Hauer, 2024; Chen et al., 2024b).
Too Low-Level Tool Usage Abstraction. Agents that require developers to interact with low-level tools increase system complexity and hinder scalability in production (Yang et al., 2024). When agents demand manual management of API calls or intricate state transitions, they become error-prone and harder to maintain. High-level abstractions, which hide tool interactions, are crucial for reducing bugs and allowing developers to focus on business logic rather than technical details.
Capability-Centric Instead of Use-Case-Centric Documentation for Tools. Documentation that focuses on tool capabilities rather than specific use cases makes it harder for FMs to leverage these tools. FMs need clear, practical examples to better operate, as seen in adaptive systems like Voyager, where the agents excel at tool use by leveraging concrete use cases. Without this, the learning curve steepens, increasing the risk of errors and slowing agent evolution (Wang et al., 2023b; Yang et al., 2024; Chen et al., 2024b). In production, use-case-driven guidance is crucial for building reliable systems (Wang et al., 2023b).
4.6. Agent Orchestration
Orchestrating multiple agents in FMware, both FM-based and traditional, ensures efficient handling of complex tasks by assigning sub-tasks to specialized agents, improving scalability and reliability (Rasal and Hauer, 2024). This mirrors microservices, where modular components collaborate for a common goal, integrating flexibility and maintainability into production systems.
Over-reliance on FM planning capability. Relying solely on an FM’s planning capabilities in production FMware introduces risks like unpredictability and lack of transparency, often leading to errors and unreliable outcomes due to issues such as hallucinations or failure to align with business logic (Guo et al., 2024a; Huang et al., 2024a). In contrast, demos may succeed with FM-based planning due to their controlled environment.
4.7. Guarding
Guarding ensures that FM and Agents using FM outputs are safe, accurate, and policy-compliant. Guardrails, for example, implements “rails” for input moderation, fact-checking, and hallucination control, maximizing reliability in production FMware (Rebedea et al., 2023; Dong et al., 2024). While demos may ignore these precautions, production systems must prioritize guardrails to ensure compliance, trust, and error prevention (Rebedea et al., 2023; Dong et al., 2024).
Simple keyword-based guarding is ineffective. Relying solely on keyword-based guarding is ineffective in production environments. Simple keyword filters can be easily bypassed or mistakenly block legitimate content due to language ambiguity. In production-ready FMware, sophisticated guardrails such as fact-checking, semantic filtering, or policy-guided decoding are essential for discerning intent and context, reducing both false positives and false negatives (Rebedea et al., 2023; Dong et al., 2024).
Lack of hallucination guardrails. Hallucinations, pose significant risks in production systems, especially in domains like healthcare or law, where factual accuracy is crucial. These fabricated outputs can lead to the spread of misinformation, undermine user trust, and even trigger legal liabilities (Ayyamperumal and Ge, 2024; Dahl et al., 2024). Additionally, FMs often lack confidence estimation, which makes it difficult to flag uncertain outputs for human review before they cause harm. User feedback systems, though essential for improving accuracy over time, are often underutilized, preventing continuous correction of inaccuracies (Ayyamperumal and Ge, 2024; Dahl et al., 2024; Spiess et al., 2024).
4.8. FMware System Testing and Optimization
FMware system testing and optimization is more complex than traditional software due to the non-deterministic nature of FMs. Unlike conventional systems, where the behavior is more predictable, FMware involves continuously evolving FMs (particularly when using third-party FMs) can silently degrade system performance without explicit changes to the underlying code.
Moreover, in production-ready FMware, system-wide testing must account for the dynamic interactions between agents, scalability concerns, and maintaining robustness under varying workloads. These challenges demand rigorous optimization, as the FMware needs to handle real-world complexities like fluctuating performance, inconsistent results, and unexpected edge cases. This goes beyond the controlled environments seen in demos, where simplicity and consistency are more achievable (Ma et al., 2024).
Lack of latency handling mechanisms. Latency is a major challenge for production-ready FMware, leading to slow responses or timeouts under real-world conditions. Ahmed et al. (2022) highlight how resource-heavy models like BERT struggle with responsiveness.  Santilli et al. (2023) show that the sequential nature of model inference introduces delays, making real-time performance difficult (Ahmed et al., 2022; Santilli et al., 2023). Variable loads further complicate the issue, as latency spikes during high demand can render systems unresponsive (Stamatescu, [n. d.]).
Lack of Retry Optimizations. Failure to implement intelligent retry strategies can significantly reduce FMware reliability, particularly when interacting with cloud-hosted FMs. In production, transient errors are common, and retry storms—where repeated requests overwhelm a system—can be costly, leading to increased error rates and decreased availability (sca, [n. d.]; RobBagby, [n. d.]). FMware also requires intelligent retries to handle scenarios like re-prompting for self-consistency, where multiple attempts may be needed to ensure the correct output (sig, [n. d.]).
High Cost of Regression Testing. Regression testing in FMware is both expensive and time-consuming due to the non-deterministic nature of FMs and the complexity of their interactions. As FM APIs evolve, performance can degrade silently, making it essential to frequently re-evaluate across numerous scenarios to ensure system reliability. This increases costs and slows down development cycles. Furthermore, traditional regression testing methods—relying on exact output matching—are inadequate for FMware, which often produce variable results. More nuanced evaluation techniques, such as similarity and correctness testing, are required to detect subtle issues in behavior. This adds to the overall complexity and cost, especially when considering large-scale applications with significant feature sets (Ma et al., 2024; sco, 2023; hum, 2023). In addition, the need for continuous evaluation across various system components in production environments demands more sophisticated testing setups, which are critical for maintaining system stability but increase both the time and resources required (Ma et al., 2024; sco, 2023).
Absence of Software Performance Engineering Practices. The absence of established software performance engineering practices, such as model swapping (replacing on FM with another to optimize memory usage), significantly complicates the optimization and scaling of FMware. In production environments, model swapping requires careful planning and compatibility testing to ensure performance gains. Without these measures, FMware risks suboptimal resource utilization, leading to increased operational costs. For instance, Li et al. (2024) and Huang et al. (2024b) stress the importance of system-level improvements in FM serving, including resource-efficient FM management to handle increasing demands. Additionally, Łazuka et al. (2024) highlight how inadequate optimization strategies can bottleneck FM inference, further reinforcing the need for robust practices.
Lack of controlled execution mechanisms. The lack of controlled execution mechanisms significantly complicates the verification of fixes and tracking the impact of updates in FMware. Without tools like feature flags, staging environments, or canary releases, testing changes before full deployment becomes risky, as updates can introduce new issues or break functionality (Wu et al., 2024a). FMware systems, like those discussed in SecGPT (Wu et al., 2024a), emphasize the importance of isolating updates and managing risks, particularly in complex environments where multiple components interact.
4.9. Deployment and Maintenance
FMware deployment requires ongoing monitoring, maintenance, and adaptation. Unlike traditional software, FMware deployment must account for non-deterministic outputs and continuous learning. In production, the focus is on ensuring reliability, efficiency, and scalability under real-world conditions.
Lack of Efficient Feedback Technology. Collecting and integrating feedback seamlessly is essential for refining models, addressing issues, and aligning with user needs. Without these mechanisms, developers miss critical insights into user behavior, leading to stagnation in performance and adaptability. Luo et al. (2024) demonstrate the importance of a robust feedback loop in the presence of continuous FM and FMware evolution, showing how the absence of efficient data collection delays improvements by failing to identify and address FM weaknesses. Production environments, unlike demos, require automated, iterative feedback systems to enable continuous learning, performance optimization, and user engagement (Wang et al., 2023a; Luo et al., 2024; aws, [n. d.]).
Lack of FMware-native observability. The lack of FMware-native observability makes it difficult to diagnose issues, optimize performance, and ensure compliance in production-ready FMware (Hassan et al., 2024b). Traditional observability tools focus primarily on functional observability metrics, such as latency, execution traces, and throughput, which are insufficient for capturing the internal reasoning and decision-making processes of FMs (Hassan et al., 2024b). This opacity hinders the ability to identify root causes of failures, especially when errors stem from complex reasoning or coordination between agents. In addition, regulatory compliance often demands detailed logs and auditing, which are difficult to maintain without comprehensive observability.
Ineffective FM Update Mechanisms Maintaining and updating FMs in production FMware presents significant issues compared to traditional software (Hassan et al., 2024b). Fixes can take months, and updates are non-deterministic—providing more training data does not ensure specific issues are resolved. Improvements may not work as expected, and previously stable features can break. In addition, improvements noted in release notes do not directly translate to features, making it difficult for developers to test what has changed or verify the improved capability. This unpredictability complicates testing and maintenance, making it difficult to prioritize efforts. In production FMware, these issues undermine reliability, performance, and user trust.
4.10. Memory Management
Memory management in FMware plays a crucial role across all stages of the lifecycle, including data alignment, prompting, grounding, agent orchestration, system optimization, and deployment. Effective memory management ensures that FMs access, store, and update relevant context and knowledge, impacting how models handle inference latency, retries, and scalability (Kwon et al., 2023; Li et al., 2024; Łazuka et al., 2024).
During deployment, memory management supports model swapping, where efficient context handling prevents performance degradation from latency issues (Łazuka et al., 2024). Additionally, memory is essential for managing retries by preserving context, preventing costly retry storms, and maintaining performance under load (RobBagby, [n. d.]; Luo et al., 2024). It also ensures that agents share state and context, crucial for coherent decision-making across agents. Robust memory management is key for optimizing dynamic workloads and ensuring regulatory compliance (Hassan et al., 2024b; Wang et al., 2023a).
Inefficient Knowledge Representation. Inefficient knowledge representation in FMware leads to storing redundant or irrelevant data, inflating computational costs and response times. For instance, Guo et al. (2024b) highlight that poor knowledge representation can cause inconsistent reasoning and incorrect outputs when retrieval systems are not optimized. In addition, the irrelevant information retrieved due to inefficient knowledge representation skew FM’s outputs (Wu et al., 2024b).
Cumbersome and Error-Prone In-Memory and Across-Memories Knowledge Management. Managing memory across multiple systems in FMware is complex and prone to errors such as synchronization issues and data corruption. Xie et al. (2023) and Chen et al. (2022) emphasize that conflicting knowledge within memory systems can lead to inconsistencies and errors when FMs must navigate between contradictory pieces of information, further complicating memory management in production environments. Packer et al. (2023) highlight that production-ready FMware requires robust memory systems, as context loss or memory mismanagement can degrade user experiences or lead to system failures (Packer et al., 2023).
4.11. Cross-cutting Issues that affect productionizing FMware
The following issues permeate across multiple stages of the FMware lifecycle and are critical challenges that we cannot ignore. Addressing these issues is essential for developing production-ready FMware that is reliable, maintainable, and of high quality.
Lack of Automated Testing Capabilities. The lack of automated testing in FMware hinders efficient issue detection and continuous integration. Existing automated tools often miss the nuanced complexities in FMware, making them inadequate for production systems where manual testing is impractical due to scale. Without automation, development slows, increasing costs and the risk of undetected regressions (Abeysinghe and Circi, 2024; Hassan et al., 2024b).
Lack of Assertion-Based Unit Tests. Writing assertion-based unit tests for FMware is difficult due to the non-deterministic nature of FMs, making expected outputs hard to define (Abeysinghe and Circi, 2024; sco, 2023). Traditional tests rely on fixed outcomes, but FMware’s varying outputs complicate this, especially in agent construction and memory management (sco, 2023). Despite these challenges, such tests are critical in production environments to detect regressions during refactoring or updates (Abeysinghe and Circi, 2024).
Text-Based Evaluation Leads to Overestimation of Quality. Text-based evaluation in FMware often leads to an overestimation of quality by failing to capture deeper issues like relevance, accuracy, or alignment with business rules. Gao et al. (2024) highlight that current text-based Natural Language Generation (NLG) evaluations often miss key aspects such as context and factual correctness. Similarly, Hasanbeig et al. (2023) argue that relying solely on surface-level outputs can cause models to overfit to specific evaluation prompts, masking underlying performance issues.
Ineffective and Inefficient AI-as-Judge Technologies. AI-as-Judge systems face inefficiency and ineffectiveness due to high costs and inherent biases. Thakur et al. (2024) highlight that models often struggle with hallucinations and inaccuracies, making their outputs unreliable in complex decision-making tasks. Additionally, the computational expense of using advanced models like GPT-4 makes them impractical for ongoing evaluations, especially when they fail to meet specific business needs (Thakur et al., 2024; Unknown, [n. d.]). These limitations, combined with the need for custom evaluation pipelines, reduce the feasibility of deploying such systems in production FMware.
Difficulty navigating the regulatory and legal compliance minefield. Legal compliance and licensing complexities permeate all stages of the FMware lifecycle. FMs must align with the legal requirements of their deployment regions; for example, using ChatGPT-4-based FMware in China may not be viable due to local regulations (Hacker et al., 2023; Kelly, [n. d.]). Licensing FM data, including synthetic data, RLHF data, and user feedback, further complicates compliance. If a model generates synthetic data, the licensing of the data it was trained on affects the licensing of the generated data. Similarly, integrating user feedback (e.g., ratings or thumbs up/down) into FMware complicates matters if users later opt out, as removing their data from the FM and related components becomes difficult. Traditional Software Bills of Materials (SBOMs) ensure compliance for conventional software, but FMware’s complexity (encompassing data, models, and FM-generated data) goes beyond what traditional SBOMs can track.
5. Challenges
Table 1. Challenges and Recurrent Issues in Developing Production-Ready FMware
Challenges Recurrent Issues
Testing Lack of assertion-based unit tests, Lack of automated testing capabilities, Text-based evaluation leads to overestimation of quality.
Observability Complexity in determining failure rationale, Lack of FMware-native observability.
Controlled Execution Lack of controlled execution mechanisms, Ineffective FM Update Mechanisms.
Resource-Aware QA Low data efficiency, Lack of latency handling mechanisms, Lack of retry optimizations, High cost of regression testing, Ineffective and inefficient AI-as-judge technologies.
Feedback Integration Lack of efficient feedback technology, Cumbersome and error-prone in-memory knowledge management.
Built-in Quality Low domain coverage, Low data quality, Lack of built-in QA checks for prompts, Non-representative or insufficient in-context learning examples, God prompts, God agents, Low information density in grounding data, Irrelevant grounding data, Over-reliance on FM planning capability, Lack of hallucination guardrails, Difficulty navigating the regulatory and legal compliance minefield.
In this section, we present the results of our semi-structured thematic synthesis and outline the key challenges in productionizing FMware. It is important to note that the challenges presented in this version of the paper address only a subset of the identified recurrent issues, as shown in Table 1. Future versions will expand on this by incorporating additional challenges that cover all the recurrent issues across the FMware lifecycle.
For each challenge, we provide an overview, a critical analysis of the state-of-the-practice, and our vision for addressing it. Tackling these challenges requires highly innovative technologies driven by cutting-edge research at the intersection of software engineering and AI. Our vision highlights these technologies, their functional requirements, and points to promising research directions.
5.1. Testing
Overview. Poor testing practices present significant roadblocks in the transition from demo FMware to production-ready systems. As identified in Section 4, the lack of robust, assertion-based unit tests, the absence of automated testing, and the over-reliance on surface-level text-based evaluation (Abeysinghe and Circi, 2024; sco, 2023; Hassan et al., 2024b) often lead to overestimation of FMware quality. In real-world production settings, they result in unreliable outputs, latent defects, and increased regression costs. To address these issues, FMware require innovative testing mechanisms that account for the unique complexities of FMs and their non-deterministic behavior.
Critical Analysis of the State-of-the-Practice. FMware testing remains an immature area, particularly in comparison to well-established practices in traditional software systems. The current testing pipeline for FMware relies heavily on manual processes for test generation. The manual identification of metamorphic relations (MRs), which are dependencies between different inputs and expected outputs, is a key bottleneck in the process. This task, while feasible for small-scale demos, becomes impractical and error-prone in production-ready FMware systems. The lack of automated test generation severely hampers the speed and coverage of testing in production systems.
Moreover, the role of AI-as-judge technologies in testing pipelines adds another layer of complexity. Current AI judges, particularly those using large FMs like GPT-4, exhibit several critical flaws in evaluating correctness. These technologies often favor outputs based on superficial attributes, such as formatting or the length of responses rather than factual accuracy or alignment with real-world constraints (Thakur et al., 2024; Chen et al., 2024a). For instance, research shows that simply reversing the order of candidate responses can dramatically alter the AI’s judgment, resulting in false positives or negatives. The failure to maintain consistency in judging introduces significant risks when scaling FMware for production environments. Research has also points to a misalignment between AI decisions and human expectations. For example, Koo et al. (Koo et al., 2024) calculate the Rank-Biased Overlap (RBO) score to measure the agreement between human preferences and model evaluations in ranking-generated texts across 16 different LLMs (RBO varies from 0 to 1 and higher values indicate higher agreement). The average RBO was 0.44, with GPT-4 scoring 0.47. Another major issue arises from the associated cost and latency with AI-as-judge systems. Relying on large models like GPT-4 is computationally expensive and not viable for long-term or large-scale production-grade testing (Parnin et al., 2023).
Our Vision. To address the limitations of current testing practices, we envision an approach that combines automated test generation and a next-generation AI-judge creator.
– Automated test generation. The process begins with gathering user feedback in real-world scenarios, such as thumbs-up or thumbs-down data, which provides rich, actionable insights. By integrating this feedback with pre-existing domain knowledge, automated metamorphic relations (MRs) are generated for specific attributes of interest. These MRs are then applied to conduct metamorphic testing (MT) on FMware, ensuring that different properties and behaviors of the system are adequately tested. Following the MT, human experts evaluate the results to identify any discrepancies or potential improvements, further refining the MRs. This iterative cycle of automated generation, testing, and human-driven feedback continuously enhances the accuracy and relevance of the MRs, ultimately leading to more robust and reliable FMware.
– Next-generation AI-judge framework. The development of such a framework is crucial for improving the reliability of evaluations. Unlike existing solutions, this framework would guide developers in crafting tailored prompts that align with specific business logic and domain constraints. These judges would be trained to focus not just on the superficial aspects of FM outputs (e.g., format, length) but on deeper attributes like factual accuracy, consistency with previous outputs, and compliance with application-specific requirements. To train the judges in a more structured, guided, flexible, evolvable, and cost-effective manner, we foresee the use of curriculum engineering (more details in Section 5.6). As a result, we would also obtain a more lightweight and efficient model, further contributing to cost-effectiveness.
5.2. Observability
Overview. Observability in FMware systems is crucial to ensure transparency, traceability, and reliability throughout their lifecycle. However, as outlined in Section 4, FMware presents unique challenges in observability due to its reliance on dynamic, non-deterministic components like FMs and autonomous agents. Recurrent issues, such as the complexity in determining the rationale behind system failures (i.e., whether stemming from limitations in prompts or FMs themselves) and the lack of FMware-native observability mechanisms (Hassan et al., 2024b), highlight the need for novel solutions. Unlike traditional software, FMware systems demand observability approaches that can capture both functional and cognitive aspects of AI agents and models.
Critical Analysis of the State-of-the-Practice. Current observability tools in FMware, while evolving, remain grounded in classical software observability practices. Tools like OpenLLMetry (tra, [n. d.]) and LangSmith (lan, [n. d.]) focus on low-level resource monitoring, capturing traces of FM calls, Vector DB interactions, and user prompts, along with metrics like latency and resource utilization. However, these frameworks overlook the cognitive processes driving FM-based decisions, particularly in complex multi-agent FMware systems where outcomes can be non-deterministic or emergent (e.g., from interactions between agents or misinterpretation of prompts).
Microsoft PowerAutomate offers one of the most advanced FMware observability frameworks by providing high-level overviews of agent workflows, enabling developers to trace where agents get stuck or fail. While an improvement, it remains focused on functional observability and does not capture how agents reason, plan, or coordinate their decisions. As FMware systems scale and become more complex, the gap between functional monitoring and decision-level observability widens.
Our Vision. Addressing the aforementioned limitations in FMware observability requires a shift in both methodology and technology. One of the key innovations needed is a general observability framework for FMware that can capture both functional performance metrics and the internal cognitive processes of FMs and agents. This framework must allow developers to probe into multiple depths of abstraction, providing traceability not only at the system level, but also across the decision-making stages of autonomous agents.
We envision an observability analytics engine that visualizes and analyzes events at higher abstraction levels, helping developers quickly pinpoint root causes in complex multi-agent workflows. To achieve this, we propose a “plane flight recorder” for agents, inspired by aviation black boxes. In FMware, this recorder would selectively capture an agent’s internal reasoning steps and communications, allowing developers to trace decision-making pathways and understand how agents reach conclusions.
However, recording an agent’s thoughts can introduce the observer effect, where observation alters behavior (e.g., adding ”think step-by-step” to a prompt changes the output). To mitigate this, we propose a surrogate agent that enables debugging without directly interfering with the system. The original problem and result are sent to the surrogate agent (whose goal is to reason rather than solve the problem), which reasons verbosely to infer the original agent’s thought process. While the surrogate’s reasoning might differ, it still provides transparency and insight into decision-making. We are exploring techniques to enhance the trustworthiness of the surrogate agent’s output.
5.3. Controlled Execution
Overview. Controlled execution in FMware development is essential for ensuring predictable, reliable, and efficient system behavior. However, as highlighted in Section 4, FMware presents a unique set of challenges due to its non-deterministic nature and the dynamic interactions between agents and models. Key recurrent issues include the lack of controlled execution mechanisms, which complicates the verification of fixes and limiting execution paths in FMware systems (Wu et al., 2024a). This lack of control severely limits debugging, reduces productivity, and can degrade system reliability, especially in multi-agent FMware where the same input might lead to divergent outputs across different executions.
Critical Analysis of the State-of-the-Practice. Traditional software engineering uses controlled execution mechanisms like feature flags and canary releases to test updates before deployment. FMware’s non-deterministic outputs and reliance on autonomous agents, however, make predicting and reproducing system behavior challenging.
A major issue is that the same input in FMware can lead to different execution paths with varying outputs due to the unpredictability of the underlying FMs and agents. Traditional techniques, which assume deterministic behavior, are not equipped for these variations, leading to flaky tests (Parnin et al., 2023). FMware often lacks repeatable execution, making it difficult to ensure consistent behavior after fixing issues. The absence of controlled execution frameworks hinders exploratory testing, preventing efficient identification of failure points and limiting performance optimizations. Without execution space restriction mechanisms like feature flags, development and maintenance become even more challenging. Additionally, the lack of detailed release notes and the disconnect between improvements and features make it hard to verify whether updates are effective. This problem is worse with cloud-hosted models, where developers have limited control over testing. Exploring and testing multiple execution paths in a structured way is essential for improving software quality and building user trust.
Our Vision. We envision a controlled execution framework designed for FMware that focuses on ensuring both repeatability and variability in execution paths, enabling comprehensive testing and validation. The framework should operate in two modes: enforcing consistent execution flows to ensure repeatability, and allowing controlled exploration of alternative flows to trigger failure points and optimize system resilience (guided exploration of the execution space).
– Repeatability. The framework should guarantee that the same input always produces the same execution flow, regardless of changes in external context or FM state. This could be achieved through execution snapshots and managing monosemantic units i.e., interpretable FM units that correspond to specific data patterns. By restoring snapshots and activating monosemantic units, developers can reproduce exact conditions that led to a bug or failure, facilitating debugging and fix verification. This approach is similar to deterministic replay systems used in distributed software, adapted for FMware’s dynamic and probabilistic nature (Alvaro and Quinn, 2024).
– Guided exploration of the execution space. This mode would systematically vary inputs, agent decisions, and model behaviors to explore different execution paths in a controlled manner. The guided process ensures comprehensive coverage, uncovering hidden bugs and allowing developers to thoroughly assess the system’s robustness across various scenarios.
5.4. Resource-Aware QA
Overview. In FMware production, managing high costs and resource demands is crucial, especially given the intensive requirements of FM operations. Traditional QA methods often overlook these needs, as FMware is hindered by low data efficiency, latency handling issues, high regression testing costs, and inefficient retry optimizations (refer Section 4). Unlike typical high-cost web services, FMs require significant computational power, have non-deterministic outputs, and need meticulous management of context and ability to quickly swap FMs. Resource-aware QA is essential for reducing the computational burden that FMware impose, particularly in large-scale environments where frequent calls to FMs can become prohibitively expensive.
Critical Analysis of the State-of-the-Practice. Current QA frameworks for FMware, adapted from traditional software engineering, fail to address FMware’s resource challenges. The problem is not just the absence of caching but the lack of FMware-native caching strategies that account for the dynamic nature of these systems. While solutions like LangChainuse traditional caching, the real challenge is deciding when to cache FM calls and ensuring cached responses remain relevant. FMware often updates databases and integrates real-time feedback, which makes cache invalidation critical. Without dynamic caching, systems risk re-running unnecessary queries or using outdated data, increasing the time and cost of regression testing.
Additionally, there is no systematic approach to optimize or prioritize the number of tests that must be run. Dependency-based test execution (Sharif et al., 2021), effective in traditional software, has not been adapted for FMware. As a result, redundant tests waste computational resources, especially where FM calls incur significant costs in terms of latency and finances.
Lastly, latency handling and retry optimizations remain underdeveloped. While traditional systems rely on retries for transient failures, FMware often needs re-prompting or alternative query structures. Current QA frameworks lack intelligent retry mechanisms that efficiently handle FM failures, leading to increased error rates and reduced system availability.
Our Vision. To address these challenges, we propose the development of a resource-aware QA framework specifically designed for FMware. This framework would incorporate a robust FMware-native caching mechanism that intelligently stores, updates and reuses FM query results, significantly reducing the need for repeated FM calls and maintaing cache integrity. By caching the results of expensive FM queries and reusing them across multiple tests, the system can reduce computational overhead while maintaining the accuracy of the testing process. This would be particularly valuable in scenarios where large volumes of tests are conducted against the same FM. In addition to caching, the framework should optimize test execution by dynamically prioritizing tests based on resource intensity. Grouping similar tests or those with expected outputs could minimize FM calls, lowering resource consumption. Test order optimization would also prioritize critical tests that are likely to surface defects, ensuring early issue detection. Native prompt compression techniques (Wingate et al., 2022) should further reduce resource demands and improve latency.
A retry optimization system is essential to handle FMware’s unpredictability. Rather than relying on basic retry logic, this system would employ dynamic prompt re-engineering and response validation. If an FM produces an unsatisfactory result, the system could modify the query intelligently, increasing the likelihood of success and preventing system overload from repeated failures.
5.5. Feedback Integration
Overview. Efficient feedback integration is vital for continuous improvement, optimization, and trustworthiness in production-ready FMware. However, as outlined in Section 4, two key issues often hinder this process: the lack of efficient feedback technology and cumbersome, error-prone memory management across different FMware systems. The absence of seamless feedback loops slows down FMware evolution and reduces adaptability, especially in complex real-world applications. Integrating feedback is further complicated by the need to manage user-specific and generalized knowledge in a scalable and non-disruptive way. As FMware becomes more widely adopted, a robust framework for feedback integration is necessary to ensure reliability, efficiency, and alignment with real-world demands.
Critical Analysis of the State-of-the-Practice. Feedback integration in FMware is underdeveloped compared to traditional software or machine learning systems, where feedback-driven optimization is well-established. The complexity of FMs and the unique characteristics of FMware complicate the application of these principles. A major issue is the lack of effective feedback mechanisms; most systems rely on ad-hoc methods requiring explicit user interactions, such as ratings (Wang et al., 2021), which suffice for demonstrations but are inadequate for production settings. Production environments demand passive, scalable, and continuous feedback collection to enable real-time insights. Without such mechanisms, developers miss essential opportunities for timely FMware refinement, hampering the correction of biases and inaccuracies and leading to slower adaptation cycles.
Moreover, current practices lack a clear framework for differentiating between types of feedback. Some feedback is universal and should be shared across all users (outer knowledge), while other feedback is specific to a single user or task (inner knowledge). Systems often lack the sophistication to manage this distinction effectively. As a result, valuable insights are overlooked, or user-specific optimizations are improperly generalized, causing inconsistencies in model performance.
Additionally, the challenge of memory management, especially across different FMware systems, compounds the difficulties of feedback integration. Managing the context and knowledge required to incorporate feedback effectively is still error-prone, with systems struggling to maintain consistency across multiple feedback loops (Peng et al., 2023). Without a robust system for synchronizing and managing feedback across different FMs and user contexts, the feedback process remains fragmented and unreliable.
Our Vision. To address these limitations, we envision a robust framework for feedback integration that builds on two key technologies: automated feedback solicitation and knowledge partitioning. These components will enable FMware systems to capture, process, and act upon feedback, ensuring continuous improvement and optimal performance.
Automated Feedback Solicitation. Automated feedback solicitation mechanisms should passively collect feedback from users during regular FMware operation, without requiring explicit input. For instance, user interaction data—such as hesitations, corrections, or patterns in query behavior—can provide valuable implicit feedback, captured and analyzed in real-time. This approach ensures that feedback is not limited to explicit ratings but includes subtle behavioral cues indicating dissatisfaction with FMware responses.
Knowledge Partitioning Framework. We propose developing a framework to manage different types of feedback by distinguishing between “outer knowledge” and “inner knowledge.” By automating the classification of feedback into these categories, FMware systems can ensure that general improvements are propagated across all users while maintaining user-specific optimizations where appropriate. The framework would also possess federated learning capabilities to leverage “inner knowledge” for global FMware improvement (Yamamoto et al., 2023). This partitioning is critical for scalable feedback integration, preventing overgeneralization of feedback while ensuring valuable information is not lost.
5.6. Built-in Quality
Overview. Built-in quality is crucial for production-ready FMware to prevent costly rework and failures. As highlighted in Section 4, recurrent issues such as low domain coverage, poor data quality, inadequate prompt validation, and lack of guardrails for hallucinations stem from an over-reliance on FMs without properly structured knowledge and action spaces. Over-reliance on FMs elicits inefficiencies and unpredictable behavior. Addressing these concerns ensures FMware systems are robust, reliable, and scalable for real-world use, not just impressive demos.
Critical Analysis of the State-of-the-Practice. Current FMware development often assumes FMs, especially LLMs, can autonomously manage knowledge and action spaces, leading to the overuse of “God prompts” and “God agents” that hinder modularization, scalability, and maintainability. Monolithic prompts complicate debugging, maintenance and scalability.
Another key issue in current practice is the lack of structured knowledge and action spaces. Many developers rely on vector databases or simple document stores that manage “raw” data without optimizing its quality or relevance, leading to low information density and degraded system performance. Current knowledge management tools do not adequately support the complexity required for real-time, high-stakes environments. Techniques like GraphRAGs improve knowledge representation but still fail to address this information density issue. Additionally, the lack of formalized curriculums for FM training results in inefficient skill acquisition and limits agents’ ability to learn compositional skills and evolve over time.
Also, efforts like SPDX 3.0 Dataset profile (Unknown, 2023) and Datasheets (Gebru et al., 2021) aimed at enabling compliance for AI-powered software fail to capture the diverse types of data in FMware such as synthetic data, and user feedback data and their compliance requirements. This omission makes it challenging to ensure legal compliance in production-ready FMware.
Finally the immature QA process for prompts and agent actions contributes to the lack of built-in quality. Most systems lack built-in prompt validation mechanisms, relying on ad-hoc testing (Kuchnik et al., 2023), which provides insufficient examples during in-context learning and leaves FMware prone to unpredictable behavior in complex tasks.
Our Vision. Our vision is to achieve high-quality FMware by effectively structuring knowledge and action spaces. This involves moving beyond vector databases to knowledge graphs that provide richer, structured information. Curriculum engineering is key to ensuring agents develop compositional skills in a structured way. Through collaborative curriculum co-creation, with AI assistance for drafting and humans-in-the-loop as reviewers, we enhance modularity, built-in quality, and skill generalization, enabling agents to build on prior knowledge rather than starting from scratch.
Developing curriculum co-creation and versioning technologies is crucial for managing the knowledge lifecycle in FMware. These tools ensure continuous evolution of knowledge bases, pruning outdated data and integrating critical updates to keep FMware systems agile and up-to-date without risking performance degradation.
Curriculum quality review and QA are essential. We propose tools to automatically optimize curricula by removing redundancies and fixing outdated or incorrect information. This ensures the data fed into FMware is relevant, dense, and actionable, improving accuracy and reliability. Built-in semantic and structural checks for prompts ensure agents receive valid, high-quality inputs, reducing errors, hallucinations, and unpredictable outputs.
Finally, for legal compliance, we propose an FMware Bill of Materials (FMwareBOM) by extending the SPDX 3.0 AI and dataset profiles (Unknown, 2023). The FMwareBOM would track all components and licenses, including synthetic data, RLHF data, and user feedback, addressing FMware’s unique complexities. A framework that automatically generates FMwareBOMs and uses formal verification techniques like SMT solvers would ensure provable compliance with legal and regulatory requirements. Integrating FMwareBOM into the development process is critical for overcoming compliance challenges and enabling the deployment of production-ready FMware.
6. Conclusion
The field of FMware is still in its early stages, and only now are organizations beginning to tackle the complexities of building production-ready FMware. Given the immense potential of FMware, the first crucial step is identifying the key challenges, so that both the research community and industry can be more aware of them and take action. Our analysis provides an initial framework for understanding these challenges, but by no means is it comprehensive. This is just the beginning—a call to the broader community to engage, share their experiences, and contribute to addressing these obstacles. We hope our paper sparks further conversations, enabling other organizations to share their own challenges and guide researchers to focus on solutions that will pave the way for establishing technologies, tools and practices to build trustworthy production-ready FMware.
Disclaimer
Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of Huawei. Also, ChatGPT with GPT-4o was used for copy-editing and table formatting. All experiments, analysis, writing, and results were performed by the authors, who also thoroughly reviewed the final content. This complies with IEEE and ACM policies on AI use in publications.